Trying to figure out this elimination problem:
3X + 5Y= 0
2X - 5Y= -25

Answer by prabhjyot(164)

3x+5y = 0
2x-5y = -25

Using Substitution Method.
Expressing x as a function of y (from the first equation) x = (0 -5y)/3
Substituting this expression into the second equation:
2(0+5y)/3-5y = -25
Simplifying this expression:
(-25)y = -75
or
y = 3
Now, substituting y = 3 back into the x expressed as a function of y:
x = (0-5y)/3 = -5